OpenStudy (anonymous):

differantation l'hopital tule ln cosx/ln cos 3x

OpenStudy (anonymous):

l 'hopital rule

OpenStudy (anonymous):

are you trying to compute a limit?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

as \(x\to 0\) ?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

ok then you can take the derivative top and bottom separately you get \[\frac{3\cos(3x)\sin(3x)}{\cos(x)\sin(x)}\] if i am not mistaken

OpenStudy (anonymous):

still get \(\frac{0}{0}\) so you have to do it again, using the product rule then you can replace \(x\) by \(0\)

OpenStudy (anonymous):

do you use quotient rule??to differentiate the equation

OpenStudy (anonymous):

L'Hopital's rule says you just take the derivative of the top, and then take the derivative of the bottom. You do not use the quotient rule.

OpenStudy (anonymous):

@satellite73 can u show me the step

OpenStudy (anonymous):

The derivative of ln cos(x) is 1/cos(x) *(-sin(x)) Similarly, the derivative of ln cos(3x) is 1/cos(3x) *(-3sin(3x)). This will give you a fraction over a fraction, which will simplify to what satellite73 posted. Unfortunately, you still end up with 0/0, so you have to use :L'Hopital's rule a second time.

OpenStudy (anonymous):

thank u @calmat01